[Hallc_running] [New Logentry] Update on Hall-C Running

latif at jlab.org latif at jlab.org
Thu Jan 25 09:40:03 EST 2018

Logentry Text:
The 1.5% C target was put back yesterday. Yesterday around 9:30 PM there was the ESR trip causing the magnets in three halls go down. Our magnets are restored. The target team will have to cool down the targets in both Hall-C and Hall-A. The cool down in Hall-A is on-going and then ours will start. Hall-B will try to finish their current data taking by the end of the day shift so that the accelerator group can do their pass change during the Hall-A/C cool down process. 

Today's day shift is canceled and swing shift will remain on standby.

We will meet today at 1:30 PM in the counting house conference room. Remote people can use the bluejeans meeting ID: 543 903 672.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3520637
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180125/66d1fcd2/attachment.html>

More information about the Hallc_running mailing list